Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Again

Expert replies
Source: — Problem Solving |

by Patrick_GMATFix » Sat Dec 21, 2013 10:13 pm
A certain store sells small, medium and large toy trucks in each of the colors red blue green and yellow. The store has an equal number of trucks of each possible color-size combination. If Paul wants a medium red truck and his mother will randomly select one of the trucks in the store, what is the probability that the truck she selects will have at least one of the two features Paul wants?

A) 1/4

B) 1/3

C) 1/2

D) 7/12

E) 2/3
The solution below is taken from the GMATFix App.

Image

-Patrick
  • Ask me about tutoring.
Join the discussion

by Stuart@KaplanGMAT » Sun Dec 22, 2013 1:31 am
A certain store sells small, medium and large toy trucks in each of the colors red blue green and yellow. The store has an equal number of trucks of each possible color-size combination. If Paul wants a medium red truck and his mother will randomly select one of the trucks in the store, what is the probability that the truck she selects will have at least one of the two features Paul wants?

A) 1/4

B) 1/3

C) 1/2

D) 7/12

E) 2/3
Hi! This question is a great illustrator of the "one minus" approach to complicated probability problems.

When you see "at least" or "at most" in probability, you're going to be solving for multiple scenarios. However, it's often quicker to solve for what you do NOT want to happen and then subtract from 1, since:

Prob(want) + Prob(don't want) = 1

and

Prob(want) = 1 - Prob(don't want)

In this case we want the truck to have at least one of two features, so what we do NOT want is a truck with neither feature.

Prob(not red) = 3/4 and Prob(not medium) = 2/3

Since we want (not red) AND (not medium) we MULTIPLY to get:

3/4 * 2/3 = 6/12 = 1/2

So, the probability of getting at least one of those two features is:

1 - 1/2 = 1/2

choose (C)!

I'm curious as to the source for this question, since it's kinder and gentler than most - usually a trap is to forget to do the final step (subtract from 1), but since the answer is 1/2 you don't get punished for forgetting!

As some good probability practice, let's calculate the long way as well:

Prob(at least 1 of red/medium) = prob(red not medium) + prob(medium not red) + prob(red AND medium)
= (1/4 * 2/3) + (3/4 * 1/3) + (1/4 * 1/3)
= 2/12 + 3/12 + 1/12
= 6/12
= 1/2

As you can see, the "one minus" approach is a lot less work!

Stuart
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by GMATGuruNY » Sun Dec 22, 2013 4:18 am
A certain store sells small, medium and large toy trucks in each of the colors red blue green and yellow. The store has an equal number of trucks of each possible color-size combination. If Paul wants a medium red truck and his mother will randomly select one of the trucks in the store, what is the probability that the truck she selects will have at least one of the two features Paul wants?

A) 1/4

B) 1/3

C) 1/2

D) 7/12

E) 2/3
To make the situation easier to evaluate, let there be ONE of each size-color combination.
Draw a CHART and put an "X" under any truck that is either red, medium or both:
Image
Of the 12 trucks in the chart, 6 are either red, medium or both.
P(selecting a good truck) = 6/12 = 1/2.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion